A merchant buys tea at the rate of Rs.4.10,Rs.3.75 and 4.50 per kg and mixes them in the proportion of 5:4:1.At what price should he sell the mixture to gain 25%?
Answer by mananth(16949)

4.10,Rs.3.75 and 4.50 per kg
suppose he mixes 5 kg,4kg & 1 kg of mix to make 10 kg
5*4.10+4*3.75+1*4.5=40
Mixture costs 40/10 = Rs.4.00 /kg
4*1.25 = Rs. 5 /kg